OpenStudy (anonymous):

simplify: 2 radical 50 - 3 radical 8

OpenStudy (anonymous):

hi\[2\sqrt{50}-3\sqrt{8}\]

OpenStudy (anonymous):

it would be \[2*5\sqrt{2}-3*2\sqrt{2}\] Because you have to have the radicals equals in order to solve

OpenStudy (anonymous):

the final answer is \[4\sqrt{2}\]